OpenStudy (anonymous):

which of them have the same result? f(2) when f(x)=3x+2 f^-1 (3) when f(x)=2x-7/3 2y+14=4y-2

OpenStudy (dumbcow):

f(2) = 3(2) +2 = 8 3 = (2x-7)/3 --> x = 8 2y+14 = 4y-2 ---> y=8 all three have same result

OpenStudy (anonymous):

are you sure??

OpenStudy (dumbcow):

yeah unless its f(x) =2x-(7/3) ??